Jorge and Anita, married taxpayers, earn $145,000 in taxable income and $55,000 in interest from an investment in City of Heflin bonds. (Use the U.S. tax rate schedule for married filing jointly). Required: If Jorge and Anita earn an additional $107,500 of taxable income, what is their marginal tax rate on this income? What is their marginal rate if, instead, they report an additional $107,500 in deductions? (For all requirements, do not round intermediate calculations. Round your answers to 2 decimal places.)

*Using 2019 Tax Rates Schedule*
Answer #1

Part A

Tax due on $145000 = 9086+(22%*(145000-78950)) = $23617

Income after addition = 145000+107500 = $252500

Tax on income after addition = 28765+(24%*(252500-168400)) = 48949

Marginal Tax Rate =(Change in Tax)/(Change in Taxable Income) = (48949-23617)/(252500-145000) = 23.56%

Part B

Tax due on $145000 = 9086+(22%*(145000-78950)) = $23617

Income after deduction = 145000-107500 = $37500

Tax on income after addition = 1940+(12%*(37500-19400)) = 4112

Marginal Tax Rate =(Change in Tax)/(Change in Taxable Income) = (4112-23617)/(37500-145000) = 18.14%